OpenStudy (anonymous):

trying to find x and y intercepts for ... y=x^2+2.2x+1.21 and y-x^2=7x

hartnn (hartnn):

hint :To get x-intercept put y=0 in the equation to get y-intercept put x=0 in the equation

OpenStudy (anonymous):

thanks i know that part lol i just dont know how to go from there...like for the first one i did y=0 then i just used the quadratic formula is that right? i got -1 and -1.2?

hartnn (hartnn):

x^2+2.2x+1.21 = 0 this can be factored, and -1 ,-1.2 is incorrect

OpenStudy (anonymous):

so i cant use the quadratic formula?

hartnn (hartnn):

those are correct, just be accurate -1,-1.21 (instead of just -1.2)

hartnn (hartnn):

and u can use formula also

OpenStudy (anonymous):

what about the second one i dont know where to start?

hartnn (hartnn):

did u get y-intercept for 1st ?

OpenStudy (anonymous):

yes i got 1.21

hartnn (hartnn):

now for 2nd eq., put x=0 for y-int., y-x^2=7x y=0 so y-intercept = (0,0) put y=0 for x-int y-x^2=7x 0=x^2+7x i believe u can solve this

hartnn (hartnn):

sorry for late reply, i got disconnected

OpenStudy (anonymous):

x=0 and x=7? right because i moved the -x^2 to the right

hartnn (hartnn):

0 and -7 ?

OpenStudy (anonymous):

its ok....and oh yea -7

hartnn (hartnn):

any more doubts ? all clear ?

OpenStudy (anonymous):

nope thank you so much :)

OpenStudy (anonymous):

can you help me with a few more though?

hartnn (hartnn):

i am already helping 2-3 ppl at same time, i'll try though...